Complete and balance each of the following equations for gas-evolution reactions.

Part A

HNO3(aq)+Na2SO3(aq)→

Express your answer as a chemical equation. Identify all of the phases in your answer.

Part B

HCl(aq)+KHCO3(aq)→

Express your answer as a chemical equation. Identify all of the phases in your answer.

Part C

HI(aq)+KHSO3(aq)→

Express your answer as a chemical equation. Identify all of the phases in your answer.

Part D

(NH4)2SO4(aq)+KOH(aq)→

Express your answer as a chemical equation. Identify all of the phases in your answer.
Answer #1

2 HNO3 (aq) + Na2SO3 (aq) = 2 NaNO3 (aq) + H2SO3 (aq)

HCl (aq) + KHCO3 (aq) = KCl (aq) + H2O (l) + CO2 (g)

HI (aq) + KHSO3 (aq) = KI (aq) + H2O (l) + SO2 (g)

(NH4)2 SO4 (aq) + 2 KOH (aq) = 2 NH4OH (aq) + H2SO4 (aq)